OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−2, 4) and a directrix of y = 6. Put the equation in standard form.

OpenStudy (anonymous):

I Just Need Answers Lol Need 2 Turn This In Like Now Teacher On My retrice

OpenStudy (anonymous):

vertex should be at \((-2,5)\) so you know it looks like \[4p(y-5)=(x+2)^2\] and all you need is \(p\)

OpenStudy (anonymous):

\(p\) is the distance from the focus to the vertex which is pretty clearly \(1\) and this parabola opens down form is \[-4(y-5)=(x+2)^2\]

OpenStudy (anonymous):

we can check it if you like

OpenStudy (anonymous):

Is That Supposed 2 Be The Answer? The -4(y-5) One

OpenStudy (anonymous):

\[-4(y-5)=(x+2)^2\]

OpenStudy (anonymous):

check it here http://www.wolframalpha.com/input/?i=parabola+-4%28y-5%29%3D%28x%2B2%29^2

OpenStudy (anonymous):

Thats Not A Answer Choice These Are... f(x) = one fourthx2 − x + 4 f(x) = −one fourthx2 − x + 4 f(x) = one fourthx2 − x + 5 f(x) = −one fourthx2 − x + 5

OpenStudy (anonymous):

then solve \[-4(y-5)=(x+2)^2\] for \(y\)

OpenStudy (anonymous):

Man Im Lost